OpenStudy (anonymous):

Cell phone companies estimate that the total number of text messages (t) since the beginning of the month can be modeled by where x is the number of days that have passed in the month. How many text messages were sent between the ninth and sixteenth day of the month? (I know that it'll end up being 40√6, but I'm stuck there)

OpenStudy (anonymous):

*modeled by t = 40√x

OpenStudy (the_fizicx99):

You sure this isn't a function?

OpenStudy (anonymous):

Absolutely, but I got the right answer already, it's 40. @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Well it's pretty easy since t(x) = \(\ 40\sqrt{x} \) \(\ t(16) = 40\sqrt{16} \rightarrow 40*4 = 160\) \(\ t(9) = 40\sqrt{9} \rightarrow 40*3 = 120\) 160 - 120 = 40

OpenStudy (the_fizicx99):

So there were 40 test messages sent between the 9th day and 16th day of that month